OpenStudy (anonymous):

I need to find the slope and y-intercept of x+5y=10

OpenStudy (anonymous):

do you know the slope intercept form of a line?

OpenStudy (anonymous):

this math confuses me, I know y=mx+b, but I am not sure how to get there from the way that the problem is stated

OpenStudy (anonymous):

okay. y = mx +b where m is the slope and b is the y intercept. now, x+5y = 10 subtract x from both sides and then divide by 5.

OpenStudy (anonymous):

x=-2

OpenStudy (anonymous):

x+5y = 10 5y = -x+10 y = -x/5 + 2 slope is -1/5, y-intercept is 2

OpenStudy (anonymous):

x+ 5y - x = 10-x subtract x from both sides 5y = -x+10 y = (-x/5) + 2 divide both sides by 5 y = (-1/5)x+2

OpenStudy (anonymous):

x+5y=10 5y= -x+10 y=-x+2 m= -1 b=2

OpenStudy (anonymous):

AntsOfDglgmut.. what would my graphing points be..

OpenStudy (anonymous):

slope is -1 and y int is 2

OpenStudy (anonymous):

well your y-intercept is just the point (0,2) you can plug in points to your original equation to find other points.

OpenStudy (anonymous):

and don't listen to kenya.

OpenStudy (anonymous):

y i did it rite

OpenStudy (anonymous):

no you didn't. you didn't divide -x by 5.